what is the mirror formula?

Grade:12th pass

4 Answers

HARI OBULESU
8 years ago
1/f=1/s+1/s’
venky
8 years ago
Mirror Formula And Magnification. In a Spherical Mirror, The distance of the object from its pole is called the Object distance (u). The distance of the Image from the Pole of the Mirror is called the Image distance (v). The Distance of the principal focus from the pole of the mirror is called the Focal length (f).
sai kumar
8 years ago
yes it is correct. 1/f = 1/s + 1/s’.
here s=object distance.
        s’=image distance.
        f=focal length.
gopathoti bujji
8 years ago
1/(focal lenth) =(1/object distance)+(1/image distance)